OpenStudy (anonymous):

Use elimination to solve the system of equations. Enter your answer as an ordered pair 2x-5y=-5 -2x+8y=-58

OpenStudy (anonymous):

is the ans (-55,-21)

OpenStudy (anonymous):

\[2x-5y=-5,~~~ -2x+8y=-58 \implies 3y=-63 \implies y=-21 \implies x=-50 \]\[\implies (x,y)=(-50,-21)\]

OpenStudy (anonymous):

can you do one more

OpenStudy (anonymous):

matiricked is correct; bad mental math by me.

OpenStudy (anonymous):

2x+6y=-8 5x-3y=88

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

mention not

OpenStudy (anonymous):

thanks matricked can you do the other problem i listed

OpenStudy (anonymous):

is the ans(14,-6)

OpenStudy (anonymous):

one more? and thnks agn

OpenStudy (anonymous):

you are a boss man 2x-y=2 3x-2y=11

OpenStudy (anonymous):

\[\left(\begin{matrix}2x+6y=-8 \\ 5x-3y=88\end{matrix}\right)\implies 12x=168 \implies x=14 \implies y=-6 \implies (x,y)=(14,-6)\]

OpenStudy (anonymous):

thanks animal

OpenStudy (anonymous):

can you one more 2x-y=2 3x-2y=11

OpenStudy (anonymous):

....

OpenStudy (anonymous):

is it (-7,-16)

OpenStudy (anonymous):

Multiply the first equation by negative two, add the equations together. That will give you the value for x, substitute to get the answer for y. Easy.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

2x+y=12 -5x-y=-33 second to last one please help t thanks agn animal and matricked

OpenStudy (anonymous):

is it (7,-2)

OpenStudy (anonymous):

5x-y=-3 15x-3y=-9 which one applys select all that do no soulotion indepedent dependent consistent infinate inconsisitent

OpenStudy (anonymous):

infinite solurtions

OpenStudy (anonymous):

Dependent, consistent. Dependent indicates infinitely many solutions. Since there is at least one solution, the system is consistent.

OpenStudy (anonymous):

thanks